# LiDDA: Data Driven Attribution at LinkedIn

### v1

@@ 2.6. Sessionization & Downsampling
[…]
2.6.1. Sessionization
First, we group impressions from certain channels in a path as a single touchpoint within a session if they are from the same campaign on the same day, as detailed in Algorithm 2. This helps reduce noise, particularly from on-platform channels where a member may observe the same campaign repeatedly due to scrolling behavior. With this approach, we model a single event however we retain the downsampling links for post-model processing to reallocate credit equally across the grouped events.
@@ 2.6.1. Sessionization
2.6.2. Downsampling
Secondly, we further address signal bias by adjusting the retention rate of specific channel, action combinations according to tunable downsampling rates. We obtain these rates using an independent analysis to estimate the comparable impact of each channel using signals such as spend and outcomes. With these ratios, we randomly drop some touchpoints to yield the expected ratio as described in Algorithm 3. We observe that this pre-processing step yields more reliable attribution results and path compositions better aligned with aggregate data.
@@ 2.6.2. Downsampling
2.7. Output Scaling
As we mentioned earlier, Media Mix Modeling (MMM) uses timeseries (daily aggregated) data to model the contribution of seasonality as well as online (display, search, video, etc.) and offline (TV, direct mail, etc.) channels. On the other hand, Data-driven Attribution (DDA) models only online channels. Online channels being independently modeled by both MMM (top-down) and DDA (bottom-up) models can lead to inconsistent results for online channels. To provide marketers a unified view for each channel performance, we apply post-modeling calibration step where DDA predictions are aligned with those from MMM models.
MMM and MTA methodologies have been used in the industry for a long time. They are often used in silos without much connection. There has been some efforts lately to connect these top-down and bottom-up approaches (Google, [n. d.]), which models each online channel contribution as a regression on aggregate channels. Another approach could be to include any prior information as a constraint in modeling (Experimental, [n. d.]). In this work, we take a different approach of scaling MTA results in post-modeling to match attribution results from MMM at the channel-level or at a different granularity.
We combine signals from our top-down MMM model with our bottom-up DDA model to create a single attribution view for the marketer. We utilize MMM to perform channel-level attribution and measure marketing incrementality in the context of macroeconomic factors, however it cannot measure specific member journeys. Conversely, DDA can perform attribution on specific interactions, but has some signal limitations. Specifically where we are unable to collect individual-level signals due to privacy restrictions or inherent channel limitations (ie. radio, events, television). Two independent attribution systems may not converge to the same point of view, particularly if they utilize different interactions. We address this by performing a post-modeling scaling across both DDA and MMM. This enables consistency at the top-level while allowing for more granular attribution within each channel to specific campaigns or engagements.
In post-modeling calibration we multiply the attribution of each touchpoint in a channel by the ratio of total number of attributed conversions by MMM to those by DDA in each channel. This does not change the relative relationship between the attribution of touchpoints from campaigns within that channel group predicted by DDA.
3. Experimental Validation and Calibration
The goal of the attribution model is to proportionally quantify the impact of each marketing action towards a conversion. Since there is no ground-truth data on the contribution of each action towards a conversion, we perform the following offline and online validations to evaluate the attribution model’s predictions.
3.1. Offline Validation
In offline validation, we seek ensure the model’s accuracy and stability as these are intuitively desirable properties of any attribution system. Because our offline process only has access to observational data, we primarily focus on typical ML performance as guardrails rather than actual attribution accuracy, as we discuss in Section 3.2.
3.1.1. Training Stability
We first assess the model’s performance using binary classification metrics, aiming for high predictive accuracy (validation ROC-AUC & PR-AUC $>0.9$ before calibration). To test the model’s sensitivity to training data, we retrain the attribution model on 10 subsets of the training data and evaluate them on the same hold out dataset. The 10 resulting models’ performance (Figure 5) and predicted action weights (Figure 6) converge, confirming that the model exhibits consistent results.
@@ 3.1.1. Training Stability
Figure 5. The distribution of ROC-AUC and PR-AUC of models retrained on 10 different subsets of training data and measured against a constant holdout. These models performance metrics converge.
Figure 6. The weight (x-axis) assigned to different actions (y-axis) in models retrained on 10 different subsets of training data (dots).
@@
3.1.2. Inference Stability
We perform a similar check on our model’s stability on inference data. For each action, we bootstrap sample the action’s weights with replacement for 100 times, calculate the average, and then analyze the distribution of the 100 averages to evaluate the stability of the predicted weights (Figure 6). We observe that the weights of each action converge. While this approach does not have an intrinsic absolute threshold, we use these values as guardrails across runs to monitor for potential issues.
Lastly, we evaluate the stability of the conversion propensities over time, using the Anderson-Darling statistic (Stephens, 1974; Anderson and Darling, 1952) on our data to evaluate whether the conversion propensities each day are from same distribution as those a week ago. We assume that the general distribution of conversion propensities should be roughly equivalent in a short period of time. The result shows that the conversion propensities each day are mostly from the same distribution.
Figure 7. Anderson-Darling statistic comparing the conversion propensities each day to those a week ago. Days where the propensities are significantly different are marked in red.
[…]
3.2. Online Validation
Offline validation does not give us insight into whether the predicted interaction weights are close to actual member behaviors. Thus, we run experiments as online validations, in which we measure the incremental effect selected campaigns have on conversions and compare it with our model predicted lift. These experiments allow us to check how close our attribution model’s predicted interaction weights align with real-world marketing efforts. We can run member-level incremental lift tests for owned channels (email and the LinkedIn platform), but we can only run geo-level experiments for external paid channels (Google Ads, Facebook, etc).
For the owned channels including email and the LinkedIn platform, we randomly split members into two groups. We select certain campaigns to hold out email sending during the experiment in the control group and conduct business as usual in the treatment group. We calculate the combined lift of the campaigns from the conversion rates of the treatment and control group. It is important to note that email sending was sparse across both members and time, limited to approximately one email per week to a small set of selected members within the connection graph. Additionally, AA tests were conducted as a basic validation of the experimental setup, and no statistically significant differences in conversion rates were found between groups, suggesting that network effects did not contaminate the results.
To validate member-level experiment, we simulate a holdout experiment within the treatment group by creating counterfactual conversion paths, where the holdout campaign touchpoints are omitted. This approach helps estimate what the conversion path would look like had the member been in the control group. This counterfactual path is a simulation of the member’s path if the customer is in the control group under two assumptions based on the experiment setup validation above: 1) members’ interaction with other campaigns were not affected by the holdout 2) members did not receive extra campaigns in place of the holdout. We use the attribution model to generate two modeled conversion probabilities:
  • $P(Conversion)$: Original path probability
  • $P^{\prime}(Conversion)$: Counterfactual path probability
We then define the lift calculation for the holdout campaign as:
(7) $Lift=\frac{P^{\prime}(Conversion)}{P(Conversion)}-1$
We compare the distribution of this simulated lift of all conversion paths against the confidence interval of the measured lift.
3.3. Calibration
There are several challenges during the experimental validation. First, the simulation of counterfactual paths by adding touches requires more assumptions such as sequencing and campaigns which may introduce noise. Second, due to the scale of our interaction data, we also do not retain the journey paths from non-converted members which complicates the measurement of control groups in model prediction. Third, the simulation can therefore only depend on the observed paths from the converted members in the treatment group which is a biased representation of outcomes for general online experiments with randomized control and treatment groups.
To mitigate the gap for selection bias in the third challenge, we fit a propensity score model for the subset of converted members in the treatment group based on the member embeddings and their group assignment, and then adjust the conversions in the test period by the inverse propensity score weighting (IPW) (Hernán and Robins, 2006). Denote $e_{i}$ as the propoensity score for member $i$ and $Z_{i}$ as the indicator for member $i$ receiving campaigns as in the treatment group (=1) and for being in the control group (=0). Denote $E_{M_{i}}$ as the member embeddings for member $i$ that is computed by LinkedIn.
Step 1: Estimate each member’s conversion propensity
(8) $e_{i}=P(Z_{i}=1|E_{M_{i}})$
Step 2: Compute the IPW
(9) $w_{i}=\begin{cases}1,&\text{if }Z_{i}=1\\ \frac{e_{i}}{1-e_{i}},&\text{if }Z_{i}=0\end{cases}$
This inverse propensity score weight can help balance the converted members in the treatment group and members in control group for fairer comparison between model predicted conversions and observed conversions from online experiments . In Figure  8, we create weight-scaled pseudo-populations through $1/w_{i}$ for those who converted in the treatment group and for those in the control group. We observe that IPW functions as expected to adjust the population from converted member in treatment group and control group to a closer match. In the next section, we apply the simulation and IPW adjustment to an online experiment result.
Figure 8. The top chart shows a mismatch between conversion propensity score distribution for converted members in the treatment group vs. control members. The bottom chart shows that conversion propensity score distributions are better matched in the pseudo-populations where we replicate the members through $1/w_{i}$ to balance the two populations.
[…]
3.4. Experiment Results
We chose three incrementality tests in the email channel to validate the attribution model and the primary metric is conversion lift defined in 7. Our goal is to evaluate the effectiveness of the acquisition email series. The workflow is illustrated in Figure 9 and the comparisons are listed in Table 2.
[…]
Test index Experimental Adjusted Simulated
1 6.67% (-2.68%, 16.02%) 9.74% (-0.66%, 20.14%) 12.54% (10.52%, 14.73%)
2 9.73% (1.71%, 17.75%) 14.05% (3.35%, 24.75%) 19.61% (16.81%, 23.44%)
3 16.65% (3.93%, 29.37%) 12.46% (-3.52%, 28.44%) 9.54% (7.53%, 11.75%)
Table 2. Experimental validation for 3 incrementality tests. Each cell contains the mean lift and its corresponding 95% confidence intervals. The confidence intervals in the first column are derived from delta method and those in the second and third columns are derived from bootstraps.
Confidence intervals for the experimental observations were derived using the delta method (van der Vaart and Wellner, 1996), and confidence intervals for the adjusted and simulated results were derived using a bootstrap with 1000 replicates. We notice that the gaps between experimental conversion lift and the simulated conversion lift were mitigated by applying the inverse propensity score weighting, and Z-score showed that adjusted conversion lift had no significant deviation from the simulated conversion lift. We concluded that our result directionally align with the experimental results.
4. Application to LinkedIn Ads Measurement
When applying LiDDA to the LinkedIn Ads use case, we also supply information about the advertiser, via a company embedding for the advertiser. This is concatenated with embeddings representing the member and their company. Similar to our GTM marketing model, we also include a touchpoint sequence representation (encoding the sequence of touch types, e.g., ”Onsite lead generation sponsored status update video ad impression”), touchpoint embeddings derived from campaign text, and positional encodings (representing the time between each touchpoint and the conversion). While GTM marketing has access to touchpoints across multiple channels, the Ads platform currently only has access to LinkedIn touchpoints. When applying the model to lead conversions obtained through LinkedIn’s Lead Generation form, on average the attribution weights show a time decay effect: touches closer to the conversion tend to get more credit (Figure 10).
Figure 10. Attribution weights for LinkedIn lead conversions show a time decay effect in aggregate; ad touches closer to the conversion tend to get more credit (i.e., higher attribution weights). Note: the attribution weights for a given conversion will always sum to one.
Meanwhile, touchpoint-level attribution from LiDDA provides very granular insights, where we see varying credit assignment to earlier touchpoints. In the provided examples in Figure 11, touches closer to the conversion tend to have higher weights; however, the weight assigned to a specific touchpoint can vary, influenced by factors like the objective type, ad format, interaction type (impression, click), ad creative, days before conversion, etc. Using these types of signals, the model can learn to allocate credit across all touchpoints in the converting paths. This data can then be surfaced to advertisers on LinkedIn’s Ad platform at different levels of aggregation, in a variety of ways, including in the Campaign Manager Reporting table (i.e., showing the number of conversions attributed to each campaign, creative, etc.), and in a plot in Measurement Insights (showing a time series of conversions attributed to different categories of campaigns).
5. Conclusion
In this paper, we introduce LiDDA, a unified transformer-based data-driven attribution approach we developed and deployed at LinkedIn. We demonstrated the innovations to DDA methodology that combines an attention mechanism, positional embeddings, touchpoint imputation, sessionization, and MMM. We also shared learnings from implementation and deployment that are broadly applicable to the Marketing and Ad tech fields.
Appendix A Additional Offline Analysis
A.1. Attention Interpretability

### v2

@@ A.2.2. Downsampling
A.3. Training Time Calibration
A.3.1. Batch‐Level Calibration
Let a training batch consist of $N$ paths. Denote by
$a^{\rm DDA}_{i}(c;\theta)=\sum_{t\in c}a_{i}^{\text{DDA}}(t;\theta),\quad(i=1,\dots,N,\;c=1,\dots,C)$
the DDA model’s attribution share for channel $c$ on path $i$. Define the batch‐aggregated channel totals
$A^{\rm DDA}_{c}(\theta)=\frac{1}{N}\sum_{i=1}^{N}a^{\rm DDA}_{i}(c;\theta),\qquad A^{\rm MMM}_{c}=a^{\rm MMM}_{c}$
where $a^{\rm MMM}_{c}$ is the global MMM‐derived share for channel $c$. We then add one of the following batch‐level penalties:
Mean‐Squared Error (MSE) Penalty
$\mathcal{L}_{\rm calib}^{\rm batch}=\sum_{c=1}^{C}\Bigl{(}A^{\rm DDA}_{c}(\theta)-A^{\rm MMM}_{c}\Bigr{)}^{2};$
Kullback–Leibler (KL) Penalty
$\mathcal{L}_{\rm calib}^{\rm batch}=\sum_{c=1}^{C}A^{\rm MMM}_{c}\,\ln\!\frac{A^{\rm MMM}_{c}}{A^{\rm DDA}_{c}(\theta)}.$
These terms are computed once per batch and combined with the base DDA loss:
$\mathcal{L}_{\rm total}=\mathcal{L}_{\rm DDA}(\theta)\;+\;\beta\,\mathcal{L}_{\rm calib}^{\rm batch}.$
Backpropagation propagates gradients from the channel‐sum residuals into each $a^{\rm DDA}_{i}(c;\theta)$.
A.3.2. Path‐Level Calibration
For each path $i$, let $\pi^{(i)}_{c}$ be the empirical fraction of touches in channel $c$:
$\pi^{(i)}_{c}=\frac{\#\{\text{touchpoints of }c\text{ in path }i\}}{\text{length}(i)}.$
Let $\pi^{\rm MMM}_{c}$ be the probability of any touchpoint’s channel being $c$ (i.e., estimate from all paths), and $a_{c}^{MMM}$ be the ”attribution” of $c$ from MMM. Define a per‐path, per‐channel weight
$w^{(i)}_{c}=\frac{\pi^{(i)}_{c}}{\pi^{\rm MMM}_{c}}.$
Form the unnormalized target
$\tilde{t}^{(i)}_{c}=w^{(i)}_{c}\;a^{\rm MMM}_{c},$
and then normalize to ”recover” a valid path-level MMM attribution
$t^{(i)}_{c}=\frac{\tilde{t}^{(i)}_{c}}{\sum_{d=1}^{C}\tilde{t}^{(i)}_{d}},\quad\sum_{c=1}^{C}t^{(i)}_{c}=1.$
We then penalize the divergence between DDA’s prediction and this target for each path:
Per‐Path MSE
$\mathcal{L}_{\rm calib}^{\rm path}=\sum_{c=1}^{C}\Bigl{(}a^{\rm DDA}_{i}(c;\theta)-t^{(i)}_{c}\Bigr{)}^{2}.$
Per‐Path KL
$\mathcal{L}_{\rm calib}^{\rm path}=\sum_{c=1}^{C}t^{(i)}_{c}\,\ln\!\frac{t^{(i)}_{c}}{a^{\rm DDA}_{i}(c;\theta)}.$
As before, the total loss is
$\mathcal{L}_{\rm total}=\mathcal{L}_{\rm DDA}(\theta)\;+\;\beta\,\mathcal{L}_{\rm calib}^{\rm path},$
and gradients flow from each path‐level residual into the DDA model parameters $\theta$.
Appendix B Additional Offline Analysis
B.1. Model Stability
We provide visualization of model stability on prediction and attribution in Figure 8, Figure 9 and Figure 3.
@@ B.1. Model Stability
Figure 8. The ROC-AUC and PR-AUC of the attributions models retrained on 10 different subsets of training data and evaluated on the same holdout dataset. The models’ performances converge.
Figure 9. Anderson-Darling statistic comparing the distribution of conversion propensities each day to the distribution from the day a week ago. Days where the propensity distributions are significantly different are marked in red.
@@
B.2. Attention Interpretability
@@ 1. Baseline
B.4. Recovering Distribution
The inverse propensity score weight can help balance the converted members in the treatment group and members in control group for fairer comparison between model predicted conversions and observed conversions from online experiments. In Figure  12, we create weight-scaled pseudo-populations through $w_{i}$ for those who converted in the treatment group and for those in the control group. We observe that IPW functions as expected to adjust the population from converted member in treatment group and control group to a closer match.
Figure 12. The top chart shows a mismatch between conversion propensity score distribution for converted members in the treatment group vs. control members. The bottom chart shows that conversion propensity score distributions are better matched in the pseudo-populations where we replicate the members through $w_{i}$ to balance the two populations.
[…]
Without IPW, we can calculate a raw estimate of $a^{\text{raw}}({\mathcal{T}^{h}})$, which can be biased.
$a^{\text{raw}}({\mathcal{T}^{h}})=\frac{\hat{P}_{1}(Y=1|S=S_{1})-\hat{P}_{0}(Y=1|S=S_{1}-\mathcal{T}^{h})}{\hat{P}_{1}(Y=1|S=S_{1})}.$
Test index $a^{\text{raw}}$ $\hat{a}^{DDA}$ $\tilde{a}^{DDA}$
1 8.72% (0.01%, 16.11%) 7.65% (4.95%, 10.07%) 12.54% (10.52%, 14.73%)
2 9.73% (1.71%, 17.75%) 14.96% (11.78%, 18.24%) 19.61% (16.81%, 23.44%)
3 16.65% (3.93%, 29.37%) 15.21% (11.65%, 19.34%) 9.54% (7.53%, 11.75%)
Table 5. Additional comparison for 3 incrementality tests. Each cell contains estimated attribution and its corresponding 95% confidence intervals.
We also calculate a plug-in estimate $\tilde{a}^{DDA}$ of the attribution with DDA model, similar as eq. [5]. Let $\pi(S)$ be the predicted conversion for touchpoint sequence $S$ from DDA model, we can compute $\tilde{P}_{1}(Y=1|S=S_{1})=(\sum_{i:Z_{i}=1}\pi(S_{1,i}))/(\sum_{i:Z_{i}=1}1)$, and $\tilde{P}_{1}(Y=1|S=S_{1}-\mathcal{T}^{h})=(\sum_{i:Z_{i}=1}\pi(S_{1,i}-\mathcal{T}^{h}))/(\sum_{i:Z_{i}=1}1)$. We compared these two estimate with ”ground truth” $\hat{a}^{\exp}$ in Table 5.
B.5. Application to Ad Platform
Touchpoint-level attribution from LiDDA provides very granular insights, where we see varying credit assignment to earlier touchpoints. In the provided examples in Figure 13, touches closer to the conversion tend to have higher weights; however, the weight assigned to a specific touchpoint can vary, influenced by factors like the objective type, ad format, interaction type (impression, click), ad creative, days before conversion, etc. Using these types of signals, the model can learn to allocate credit across all touchpoints in the converting paths. This data can then be surfaced to advertisers on LinkedIn’s Ad platform at different levels of aggregation, in a variety of ways, including in the Campaign Manager Reporting table (i.e., showing the number of conversions attributed to each campaign, creative, etc.), and in a plot in Measurement Insights (showing a time series of conversions attributed to different categories of campaigns).